# InfraFix: Technology-Agnostic Repair of Infrastructure as Code

### v1

@@ Abstract.
Infrastructure as Code (IaC) enables scalable and automated IT infrastructure management but is prone to errors that can lead to security vulnerabilities, outages, and data loss. While prior research has focused on detecting IaC issues, Automated Program Repair (APR) remains underexplored, largely due to the lack of suitable specifications. In this work, we propose InfraFix, the first technology-agnostic framework for repairing IaC scripts. Unlike prior approaches, InfraFix allows APR techniques to be guided by diverse information sources. Additionally, we introduce a novel approach for generating repair scenarios, enabling large-scale evaluation of APR techniques for IaC. We implement and evaluate InfraFix using an SMT-based repair module and a state inference module that uses system calls, demonstrating its effectiveness across 254,755 repair scenarios with a success rate of 95.5%. Our work provides a foundation for advancing APR in IaC by enabling researchers to experiment with new state inference and repair techniques using InfraFix and to evaluate their approaches at scale with our repair scenario generation method.
@@ 1. Introduction
Given this, it is crucial to understand how we can detect and prevent errors in IaC. Currently, the IaC research community is primarily focused on error detection (Saavedra and Ferreira, 2022; Saavedra et al., 2023; Rahman et al., 2021; Reis et al., 2022; Hassan et al., 2024; Begoug et al., 2024; Sotiropoulos et al., 2020; Opdebeeck et al., 2022; Rahman et al., 2023; Lepiller et al., 2021; Schwarz et al., 2018; Sharma et al., 2016; Bessghaier et al., 2024; Opdebeeck et al., 2023; Zerouali et al., 2023). Only a few studies have taken the additional step of exploring Automated Program Repair (APR) in IaC (Weiss et al., 2017; Hassan et al., 2018), and most of these efforts focus on static repair rules that do not depend on program-specific specifications (Bui et al., 2023; Henkel et al., 2021; Durieux, 2024). Since most APR techniques use weak specifications, such as test suites (Goues et al., 2019; Monperrus, 2018b, a), we believe that the development of APR in IaC is being hindered by the lack of adoption of testing practices for IaC, a problem identified in the literature (Guerriero et al., 2019; Sokolowski et al., 2024a; Rahman et al., 2019a; Sokolowski et al., 2024b).
@@
We make the following contributions.
We propose InfraFix, the first technology-agnostic framework that can automatically repair IaC scripts. We provide an implementation that includes an example of an SMT-based repair module and a state inference module.
We introduce a novel approach for the generation of repair scenarios that can be used to evaluate APR approaches for IaC. We use this approach to evaluate our implementation of InfraFix in 254,755 different repair scenarios, achieving a successful repair rate of 95.5%.
Tool Availability
A replication package that includes all the datasets of IaC scripts used to create repair scenarios, the scripts used to run our generation approach, and the source code of the implementation of InfraFix is available at: https://doi.org/10.5281/zenodo.15061764. A video describing and demonstrating InfraFix is available at: https://youtu.be/ETSW0n9DcdU. InfraFix is open-source and available at: https://github.com/sr-lab/GLITCH/tree/interactive_repair.
@@ 2.1. State Inference Module
A state inference module is responsible for creating the specification that guides the repair of the IaC script. This module receives as input information about the system and a list of resources affected by that information and infers the desired system state. A system state is a list of resource states. A resource state consists of the resource’s identifier and a dictionary with the attributes of the resource. Due to the declarative nature of IaC scripts, system states are a suitable specification to guide APR techniques.
We can use different sources of information to infer states, such as system calls from executed shell commands (Weiss et al., 2017), cloud logs such as AWS CloudTrail (Amazon Web Services, Inc., 2025) or Google Cloud Audit Logs (Google, 2025), GitHub issues, or other relevant data.
The current implementation of InfraFix extends Weiss et al.’s approach, providing a state inference module that uses system calls to get a list of files affected by a sysadmin’s shell commands. The module then uses that list to retrieve the attributes of those files from the file system.
Importantly, InfraFix is flexible and can be easily extended with other valuable approaches. For example, consider the scenario where a sysadmin is asked to install Steam111Steam is a video game digital distribution service and storefront managed by Valve (Wikipedia, 2024). on all Linux machines in a gaming convention. The sysadmin develops an Ansible script that installs the steam package on these machines. However, when trying to deploy the script, it returns the following error:
@@ 2.2. Technology-Agnostic Approach
To ensure InfraFix is technology-agnostic, we need to address two main challenges:
Syntactic structure: Different IaC technologies may use different configuration languages. Dealing simultaneously with different configuration languages requires an abstraction that overcomes the different syntactic structures.
Naming: Depending on the IaC technology, attributes, reserved values, or resource types with the same semantic meaning may have different names. For instance, in Chef, links and directories are managed using distinct resources, link and directory, whereas in Ansible and Puppet, files, links, and directories are all handled within the same file module.
@@ 2.2.2. Normalization Database
One might argue that only standard system resources share enough overlap across IaC technologies to justify normalization. However, since these technologies configure the same types of resources, configuration requirements naturally create commonalities, making the normalization process broadly applicable. For example, the community.aws.lightsail module in Ansible (Ansible project contributors, 2025) overlaps significantly with the aws_lightsail_instance resource in Terraform (HashiCorp, 2025).
@@ 2.3. Repair Module
For this paper, we integrate the SMT-based approach of Weiss et al. (Weiss et al., 2017) into InfraFix, but we extend it to support repairs for other types of resources other than files. Currently, this module supports the repair of the following resources: files, packages, services and users in Ansible, Puppet, and Chef; AWS IAM roles, AWS EC2 instances, and AWS S3 buckets in Terraform and Ansible. Furthermore, we improved the approach of Weiss et al. by incorporating support for conditional branching, attribute sketching (Solar-Lezama et al., 2006), and direct handling of attributes without requiring preprocessing of IaC scripts, thus expanding the spectrum of potential repairs.
@@ 3.1. Datasets
To evaluate InfraFix, we use the Ansible and Chef datasets provided by Saavedra and Ferreira (Saavedra and Ferreira, 2022), the Puppet datasets provided by Rahman et al. (Rahman et al., 2019b), and the Terraform dataset provided by Gonçalves (Gonçalves, 2023). We also included in our evaluation the Puppet files used to evaluate Tortoise (Weiss et al., 2017). Subsequently, we curated the IaC scripts in the selected datasets based on the following criteria: (1) Test or library scripts are discarded. (2) If two or more files contain the same exact content, only one is kept. (3) Files that are not parsable by InfraFix due to errors are discarded. After applying all the criteria, we obtain five curated datasets. To obtain our evaluation benchmarks, we filter out all files that do not contain at least one resource currently supported by InfraFix. Table 1 shows the number of scripts in the datasets after each step.
@@ 3.2. Automated Generation of Repair Scenarios
First, we transform the IaC script into the normalized GLITCH’s IR. Then, we employ a state inference module that takes the IR as input to determine the possible system state(s) resulting from applying the original IaC script. We note that multiple potential system states may arise when it is not possible to statically determine which branch of a conditional statement will be executed. After obtaining the possible system states, we apply a sequence of mutations to each state, obtaining a set of desired system states. Mutations involve modifying the attributes of the resources in the state of the system with arbitrary yet valid values.
Each desired system state corresponds to a repair scenario. For each scenario, we execute InfraFix with the normalized IR and the desired system state. Since InfraFix can generate multiple possible solutions, we consider the repair successful if any of the fixed GLITCH’s IR results in a system state equivalent to the desired one. If InfraFix raises an error, we consider the state of the repair scenario to be erroneous. If there is a timeout without finding any solutions, we consider the repair scenario to be timed out. Otherwise, we consider it failed. The timeout is set to 120 seconds.
@@ 3.3. Results
Tortoise | 1,161 | 1,061 (91.4%) | 80 (6.89 %) | 0 (0.0%) | 20 (1.7%) |
Total | 254,755 | 243,404 (95.5%) | 2,318 (1.0%) | 219 (0.0%) | 8,347 (3.3 %) |
We successfully generated repair scenarios for 7,833/12,358 (Ansible), 125/4,868 (Puppet), 2,069/3,568 (Chef), 736/1,300 (Terraform), and 13/13 (Tortoise) files. From these files, we generated 183,825 repair scenarios for Ansible, 4,868 for Puppet, 61,357 for Chef, 3,544 for Terraform, and 1,161 for Tortoise. The experiments were conducted on a machine with an Intel Xeon Silver 4110 CPU @ 2.10GHz and 64GB of RAM. We used 32 cores to parallelize the evaluation with all repair scenarios of a single file being handled by the same process. Table 2 shows the results obtained for each benchmark.
InfraFix passed 95.8%, 79.5%, 96.5%, 100.0%, and 91.4% of the repair scenarios in the Ansible, Puppet, Chef, Terraform, and Tortoise benchmarks, respectively. Upon manually examining the repair scenarios, we found most failures were due to: (1) references to undefined variables assigned to attributes or other variables, (2) attributes containing shared variable references that, when modified, unintentionally affect multiple attributes, and (3) defining multiple resources within the same construct using an array as title (Puppet, Inc., 2025).
@@ 4. Related Work
Due to space constraints, we focus on repair techniques specific to IaC. Bui et al. (Bui et al., 2023) and Durieux et al. (Durieux, 2024) developed tools for automated detection and repair of code smells in Dockerfiles. These works are similar to ours in their attempt to repair IaC scripts. However, they focus on code smells, while we consider a wider range of issues. Henkel et al. used machine learning to infer repair rules based on build log analysis. They developed a tool, called Shipwright, with the inferred rules to repair Dockerfiles that fail to build (Henkel et al., 2021). The main difference between Shipwright and InfraFix is that Shipwright relies on inferred and pre-defined rules, while our technique uses information to infer the system state and guide the repair. Hassan et al. proposed RUDSEA, an approach that automatically recommends updates to Dockerfiles based on changes in the source code and build configuration files (Hassan et al., 2018). Weiss et al. (Weiss et al., 2017) proposed Tortoise, a tool that implements a technique called imperative configuration repair, which enables automatic repair of IaC scripts guided by a series of shell commands. Similarly to RUDSEA, Tortoise is focused on a single IaC technology. Our approach is more general, both in terms of resources and IaC technologies supported. In fact, the specifications used by RUDSEA to guide repair could be used with InfraFix, provided that a state inference mechanism is implemented.
@@ Acknowledgements.
We would like to thank Margarida Ferreira for her help with string encoding, Aaron Weiss and Arjun Guha for their assistance with Tortoise, Vasco Manquinho and Inês Lynce for their feedback, and Bruno Saavedra for his help with the video. This work was supported by FCT, Fundação para a Ciência e a Tecnologia under grant BD/04736/2023, and projects 2024.07411.IACDC and UIDB/50021/2020 (DOI: 10.54499/UIDB/50021/2020).References
@@
- Amazon Web Services, Inc. (2025) Amazon Web Services, Inc. 2025. What is AWS CloudTrail? https://docs.aws.amazon.com/awscloudtrail/latest/userguide/cloudtrail-user-guide.html [Accessed 21-03-2025].
@@
- Bui et al. (2023) Quang-Cuong Bui, Malte Laukötter, and Riccardo Scandariato. 2023. DockerCleaner: Automatic Repair of Security Smells in Dockerfiles. In 2023 IEEE International Conference on Software Maintenance and Evolution (ICSME). IEEE, 160–170.
- Durieux (2024) Thomas Durieux. 2024. Empirical study of the docker smells impact on the image size. In Proceedings of the IEEE/ACM 46th International Conference on Software Engineering. 1–12.
@@
- Google (2025) Google. 2025. Cloud Audit Logs overview. https://cloud.google.com/logging/docs/audit [Accessed 21-03-2025].
@@
- Guerriero et al. (2019) Michele Guerriero, Martin Garriga, Damian A Tamburri, and Fabio Palomba. 2019. Adoption, support, and challenges of infrastructure-as-code: Insights from industry. In 2019 IEEE international conference on software maintenance and evolution (ICSME). IEEE, 580–589.
@@
- Janardhan (2021) Santosh Janardhan. 2021. More details about the October 4 outage. https://engineering.fb.com/2021/10/05/networking-traffic/outage-details/ [Accessed 25-01-2024].
@@
- Sokolowski et al. (2024a) Daniel Sokolowski, David Spielmann, and Guido Salvaneschi. 2024a. Automated Infrastructure as Code Program Testing. IEEE Transactions on Software Engineering (2024).
@@
- Solar-Lezama et al. (2006) Armando Solar-Lezama, Liviu Tancau, Rastislav Bodik, Sanjit Seshia, and Vijay Saraswat. 2006. Combinatorial sketching for finite programs. In Proceedings of the 12th International Conference on Architectural Support for Programming Languages and Operating Systems. 404–415.
@@
- Wikipedia (2024) Wikipedia. 2024. Steam (service) — Wikipedia, The Free Encyclopedia. http://en.wikipedia.org/w/index.php?title=Steam%20(service)&oldid=1244633065. [Accessed 09-09-2024].

### v2

@@ Abstract.
Infrastructure as Code (IaC) enables scalable and automated IT infrastructure management but is prone to errors that can lead to security vulnerabilities, outages, and data loss. While prior research has focused on detecting IaC issues, Automated Program Repair (APR) remains underexplored, largely due to the lack of suitable specifications. In this work, we propose InfraFix, the first technology-agnostic framework for repairing IaC scripts. Unlike prior approaches, InfraFix allows APR techniques to be guided by diverse information sources. Additionally, we introduce a novel approach for generating repair scenarios, enabling large-scale evaluation of APR techniques for IaC. We implement and evaluate InfraFix using an SMT-based repair module and a state inference module that uses system calls, demonstrating its effectiveness across 254,288 repair scenarios with a success rate of 95.7%. Our work provides a foundation for advancing APR in IaC by enabling researchers to experiment with new state inference and repair techniques using InfraFix and to evaluate their approaches at scale with our repair scenario generation method.
@@ 1. Introduction
Given this, it is crucial to understand how we can detect and prevent errors in IaC. Currently, the IaC research community is primarily focused on error detection (Saavedra and Ferreira, 2022; Saavedra et al., 2023; Rahman et al., 2021; Reis et al., 2022; Hassan et al., 2024; Begoug et al., 2024; Sotiropoulos et al., 2020; Opdebeeck et al., 2022; Rahman et al., 2023; Lepiller et al., 2021; Schwarz et al., 2018; Sharma et al., 2016; Bessghaier et al., 2024; Opdebeeck et al., 2023; Zerouali et al., 2023). Only a few studies have taken the additional step of exploring Automated Program Repair (APR) in IaC (Weiss et al., 2017; Hassan et al., 2018), and most of these efforts focus on static repair rules that do not depend on program-specific specifications (Bui et al., 2023; Henkel et al., 2021; Durieux, 2024). Since most APR techniques use weak specifications, such as test suites (Goues et al., 2019; Monperrus, 2018b, a), the development of APR in IaC is being hindered by the lack of adoption of testing practices for IaC, as noted in prior work (Guerriero et al., 2019; Sokolowski et al., 2024a; Rahman et al., 2019a; Sokolowski et al., 2024b).
@@
We make the following contributions: (1) We propose InfraFix, the first technology-agnostic framework that can automatically repair IaC scripts. Our implementation includes an example of an SMT-based repair module and a state inference module. (2) We introduce a novel approach for the generation of repair scenarios that can be used to evaluate APR approaches for IaC. We use this approach to evaluate our implementation of InfraFix in 254,288 different repair scenarios, achieving a successful repair rate of 95.7%.
Tool Availability. A replication package with all the datasets of IaC scripts used to create repair scenarios, the scripts used to run our generation approach, and the source code of the implementation of InfraFix is available at: https://doi.org/10.5281/zenodo.15061764. A video describing and demonstrating InfraFix is available at: https://youtu.be/ETSW0n9DcdU. InfraFix is open-source and available at: https://github.com/sr-lab/GLITCH/tree/issta2025.
@@ 2.1. State Inference Module
A state inference module creates the specification that guides the repair of the IaC script. This module receives as input information about the system, such as logs or system calls, and a list of resources affected by that information and infers the desired system state. A system state is a list of resource states. A resource state consists of the resource’s identifier and a dictionary with the attributes of the resource. Due to the declarative nature of IaC scripts, system states are a suitable specification to guide APR techniques. We can use different sources of information to infer states, such as system calls from executed shell commands (Weiss et al., 2017), cloud logs such as AWS CloudTrail or Google Cloud Audit Logs, or GitHub issues.
The current implementation of InfraFix extends Weiss et al.’s approach, providing a state inference module that uses system calls to get a list of files affected by a sysadmin’s shell commands. The module then uses that list to retrieve the attributes of those files from the file system. Importantly, InfraFix is flexible and can be easily extended with other valuable approaches. For example, consider the scenario where a sysadmin is asked to install Steam111Steam is a video game digital distribution service managed by Valve. on all Linux machines in a gaming convention. The sysadmin develops an Ansible script that installs the steam package on these machines. However, when trying to deploy the script, it returns the following error:
@@ 2.2. Technology-Agnostic Approach
To ensure InfraFix is technology-agnostic, we need to address two main challenges: (1) Syntactic structure:Different IaC technologies may use different configuration languages. Dealing simultaneously with different configuration languages requires an abstraction that overcomes the different syntactic structures. (2) Naming:Depending on the IaC technology, attributes, reserved values, or resource types with the same semantic meaning may have different names. For instance, in Chef, links and directories are managed using distinct resources, whereas in Ansible and Puppet, files, links, and directories are all handled within the same file module.
@@ 2.2.2. Normalization Database
While only standard system resources may seem to justify normalization, common configuration needs across IaC tools create natural overlaps, making the normalization process broadly applicable. For example, the community.aws.lightsail module in Ansible (Ansible project contributors, 2025) overlaps significantly with the aws_lightsail_instance resource in Terraform (HashiCorp, 2025).
@@ 2.3. Repair Module
We integrate the SMT-based approach of Weiss et al. (Weiss et al., 2017) into InfraFix, but we extend it to support repairs for other types of resources other than files. We adopt this approach as it represents the state-of-the-art in APR for IaC. Currently, this module supports the repair of the following resources: files, packages, services and users in Ansible, Puppet, and Chef; AWS IAM roles, AWS EC2 instances, and AWS S3 buckets in Terraform and Ansible. We selected these types of resources because they are shared across the supported IaC technologies and rank among the most frequently used, as evidenced by the statistics provided in our replication package. Furthermore, we improved the approach of Weiss et al. by incorporating support for conditional branching, attribute sketching (Solar-Lezama et al., 2006), and direct handling of attributes without requiring preprocessing of IaC scripts, thus expanding the spectrum of potential repairs.
@@ 3.1. Datasets
To evaluate InfraFix, we use the Ansible and Chef datasets provided by Saavedra and Ferreira (Saavedra and Ferreira, 2022), the Puppet datasets provided by Rahman et al. (Rahman et al., 2019b), and the Terraform dataset provided by Gonçalves (Gonçalves, 2023). We use these datasets since they have been previously used for large-scale empirical studies on IaC. We also included in our evaluation the Puppet files used to evaluate Tortoise (Weiss et al., 2017). Subsequently, we curated the IaC scripts in the selected datasets based on the following criteria: (1) Test or library scripts are discarded. (2) If two files contain the same content, only one is kept. (3) Files that are not parsable by InfraFix due to errors are discarded. After applying all the criteria, we obtain five curated datasets. To obtain our evaluation benchmarks, we filter out all files that do not contain at least one resource currently supported by InfraFix. Table 1 shows the number of scripts in the datasets after each step.
@@ 3.2. Automated Generation of Repair Scenarios
First, we transform the IaC script into the normalized GLITCH’s IR. Then, we employ a state inference module that takes the IR as input to determine the possible system state(s) resulting from applying the original IaC script. We note that multiple potential system states may arise when it is not possible to statically determine which branch of a conditional statement will be executed. After obtaining the possible system states, we apply a sequence of mutations to each state, obtaining a set of desired system states. These mutations simulate potentially desired modifications to the system state. Mutations involve modifying the attributes of the resources in the state of the system with arbitrary yet valid values.
Each desired system state corresponds to a repair scenario. For each scenario, we execute InfraFix with the normalized IR and the desired system state. Since InfraFix can generate multiple possible solutions, we consider the repair successful if any of the fixed GLITCH’s IR results in a system state equivalent to the desired one. If InfraFix raises an error, the scenario is marked as erroneous. If there is a timeout without finding any solutions, we consider the repair scenario to be timed out. Otherwise, we consider it failed. The timeout threshold is set to 120 seconds.
@@ 3.3. Results
Tortoise | 1,161 | 1,064 (91.7%) | 80 (6.9 %) | 0 (0.0%) | 17 (1.5%) |
Total | 254,288 | 243,407 (95.7%) | 2,318 (0.9%) | 219 (0.1%) | 8,344 (3.3 %) |
We successfully generated repair scenarios for 7,833/12,358 (Ansible), 125/4,868 (Puppet), 2,069/3,568 (Chef), 736/1,300 (Terraform), and 13/13 (Tortoise) files. Failures were due to two main reasons: (1) although these files contained supported resources, their names could not be statically resolved to strings; and (2) the resources were defined within other constructs, such as function calls or classes, that were not invoked or instantiated within the same script. As a result, these resources could not be included in the inferred system state for those files.
We generated 183,825 repair scenarios for Ansible, 4,868 for Puppet, 61,357 for Chef, 3,544 for Terraform, and 1,161 for Tortoise. The experiments were conducted on a machine with an Intel Xeon Silver 4110 CPU @ 2.10GHz and 64GB of RAM. We used 32 cores to parallelize the evaluation with all repair scenarios of a single file being handled by the same process. Table 2 shows the results obtained for each benchmark. InfraFix passed 95.8%, 79.5%, 96.5%, 100.0%, and 91.4% of the repair scenarios in the Ansible, Puppet, Chef, Terraform, and Tortoise benchmarks, respectively. Upon manually examining the repair scenarios, we found most failures were due to: (1) references to undefined variables assigned to attributes or other variables, (2) attributes containing shared variable references that, when modified, unintentionally affect multiple attributes, and (3) defining multiple resources within the same construct using an array as title (Puppet, Inc., 2025).
@@ 4. Related Work
Due to space constraints, we focus on repair techniques specific to IaC. Bui et al. (Bui et al., 2023) and Durieux et al. (Durieux, 2024) developed tools for automated detection and repair of code smells in Dockerfiles. These works are similar to ours in their attempt to repair IaC scripts. However, they focus on code smells, while we consider a wider range of issues. Henkel et al. used machine learning to infer repair rules based on build log analysis. They developed a tool, called Shipwright, with the inferred rules to repair Dockerfiles that fail to build (Henkel et al., 2021). The main difference is that Shipwright uses pre-defined and inferred rules, while InfraFix infers system state to guide repairs. Hassan et al. proposed RUDSEA, an approach that automatically recommends updates to Dockerfiles based on changes in the source code and build configuration files (Hassan et al., 2018). Weiss et al. proposed Tortoise, a tool that implements a technique called imperative configuration repair, which enables automatic repair of IaC scripts guided by a series of shell commands (Weiss et al., 2017). Similarly to RUDSEA, Tortoise is focused on a single IaC technology. Our approach is more general, both in terms of resources and IaC technologies supported. In fact, the specifications used by RUDSEA could be used with InfraFix, provided that a state inference module is implemented.
@@ Acknowledgements.
We would like to thank Margarida Ferreira for her help with string encoding, Aaron Weiss and Arjun Guha for their assistance with Tortoise, Vasco Manquinho and Inês Lynce for their feedback, and Bruno Saavedra for his help with the video. This work was supported by Fundação para a Ciência e a Tecnologia (FCT): N. Saavedra by grant BD/04736/2023; N. Saavedra and J. F. Ferreira by projects UIDB/50021/2020 (DOI: 10.54499/UIDB/50021/2020) and the ‘InfraGov’ project, with ref. n. 2024.07411.IACDC (DOI: 10.54499/2024.07411.IACDC), funded by the ‘Plano de Recuperação e Resiliência (PRR)’ under the investment ‘RE-C05-i08 - Ciência Mais Digital’, measure ‘RE-C05-i08.M04’ (in accordance with the FCT Notice No. 04/C05 i08/2024), framed within the financing agreement signed between the ‘Estrutura de Missão Recuperar Portugal (EMRP)’ and the FCT as an intermediary beneficiary. A. Mendes was financed by National Funds through the Portuguese funding agency, FCT, within project LA/P/0063/2020 (DOI: 10.54499/LA/P/0063/2020).References
@@
- Bui et al. (2023) Quang-Cuong Bui, Malte Laukötter, and Riccardo Scandariato. 2023. DockerCleaner: Automatic Repair of Security Smells in Dockerfiles. In 2023 IEEE International Conference on Software Maintenance and Evolution. IEEE, 160–170.
- Durieux (2024) Thomas Durieux. 2024. Empirical study of the docker smells impact on the image size. In Proceedings of the IEEE/ACM 46th Int. Conf. on Software Engineering. 1–12.
@@
- Guerriero et al. (2019) Michele Guerriero, Martin Garriga, Damian A Tamburri, and Fabio Palomba. 2019. Adoption, support, and challenges of infrastructure-as-code: Insights from industry. In 2019 IEEE Int. Conf. on Software Maintenance and Evolution. 580–589.
@@
- Janardhan (2021) Santosh Janardhan. 2021. More details about the October 4 outage. https://engineering.fb.com/2021/10/05/networking-traffic/outage-details/
@@
- Sokolowski et al. (2024a) Daniel Sokolowski, David Spielmann, and Guido Salvaneschi. 2024a. Automated Infrastructure as Code Program Testing. IEEE TSE (2024).
@@
- Solar-Lezama et al. (2006) Armando Solar-Lezama, Liviu Tancau, Rastislav Bodik, Sanjit Seshia, and Vijay Saraswat. 2006. Combinatorial sketching for finite programs. In ASPLOS. 404–415.